# GridRoute: A Benchmark for LLM-Based Route Planning with Cardinal Movement in Grid Environments

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) have demonstrated their potential in planning and reasoning tasks, offering a flexible alternative to classical pathfinding algorithms. However, most existing studies focus on LLMs’ independent reasoning capabilities and overlook the potential synergy between LLMs and traditional algorithms. To fill this gap, we propose a comprehensive evaluation benchmark GridRoute to assess how LLMs can take advantage of traditional algorithms. We also propose a novel hybrid prompting technique called Algorithm of Thought (AoT), which introduces traditional algorithms’ guidance into prompting. Our benchmark evaluates six LLMs ranging from 7B to 72B parameters across various map sizes, assessing their performance in correctness, optimality, and efficiency in grid environments with varying sizes. Our results show that AoT significantly boosts performance across all model sizes, particularly in larger or more complex environments, suggesting a promising approach to addressing path planning challenges.111Our code is open-sourced at https://github.com/LinChance/GridRoute.
@@
Kechen Li1, Yaotian Tao2, Ximing Wen3, Quanwei Sun2, Zifei Gong4, Chang Xu2, Xizhe Zhang2, Tianbo Ji1,††thanks: Corresponding author 1School of Transportation and Civil Engineering, Nantong University, China 2HongKong Linchance Intelligent Technology Co., Ltd. (LinChance), China 3College of Computing and Informatics, Drexel University, USA 4University of New South Wales, Australia kechen.li@stmail.ntu.edu.cn, {taoyaotian,quanweisun,xuchang,zhangxizhe}@linchance.com xw384@drexel.edu, z3224950@zmail.unsw.edu.au, jitianbo@ntu.edu.cn
@@ 1 Introduction
To bridge this gap, we propose a comprehensive benchmark GridRoute to systematically evaluate the role of traditional algorithmic guidance in enhancing the pathfinding capabilities of LLMs. Our benchmark is designed around three core prompting paradigms: (1) independent prompting, (2) Algorithm of Thought(AoT) prompt, where traditional algorithms such as DFS, A*, and Dijkstra provide reasoning guidance; and (3) hybrid prompting, which leverages algorithm-generated trajectories or decisions within both direct-answer and reasoning-based prompts. This design facilitates a comprehensive comparison of traditional algorithmic guidance versus independent prompting approaches, and further allows for in-depth analysis of variation within algorithm-guided prompts.
@@
Through this benchmark, we aim to answer several key questions: To what extent can LLMs benefit from classical search heuristics? How do the different shot strategies affect the traditional algorithm’s guidance for LLM? Do algorithm-guided prompt performance remain indispensable regardless of model scale and map size? Our findings suggest that algorithm-guided prompts achieve better results than vanilla prompts and comparable performance compared with the CoT prompt. While larger models tend to perform better under independent prompting, algorithm-guided strategies significantly boost performance across all model sizes, particularly in larger or more complex environments. Moreover, integrating few-shot learning could further improve the performance of algorithm algorithm-guided prompt. These results highlight the complementary strengths of LLMs and classical algorithms and point toward promising directions for hybrid neuro-symbolic planning systems.
@@
We introduce AoT prompting, a novel framework that guides LLMs using classical pathfinding algorithms (e.g., A*, Dijkstra, DFS) by embedding their reasoning principles directly into the prompt.
We perform large-scale experiments across multiple LLMs and environment sizes, showing that AoT consistently improves planning performance, especially in complex settings, and offers complementary benefits to traditional CoT prompting.
@@ 4.2.1 Independent Route Planning Prompt
4.2.2 Algorithms of Thought (AoT) Prompt
To evaluate the capability of LLMs in applying classical algorithms, we designed the Algorithms of Thought (AoT) prompting framework. This framework guides LLMs to mimic classical path planning algorithms by prompting them to apply algorithmic reasoning, analyze the task environment, constrain movement directions, explain algorithmic principles, and adhere to a clear output format.
AoT-DFS: This prompt is designed to guide the model in following the Depth-First Search (DFS).Instead of outlining the algorithm formally, the prompt incorporates its core mechanisms—ordered expansion, backtracking, and visit marking—into the structure of the instruction itself. By structuring the order of exploration and discouraging revisits,it leads the model to follow a behavior pattern similar to DFS.
AoT-A*: This prompt adapts the logic of the A* algorithm to guide the model in path planning. It introduces key elements of A*, such as cost estimation, heuristic evaluation, and goal-directed prioritization. The prompt defines valid moves, explains how to measure distance using the Manhattan heuristic, and describes how to compare path options. The model is prompted to evaluate each possible move based on its distance to the goal and the cost incurred so far, simulating how A* selects efficient paths.
AoT-Dijkstra: It provides a structured framework by embedding Dijkstra’s core logic – greedy distance minimization, neighbor evaluation, and path optimality. In detail, it enforces a multi-step workflow: initializing Dijkstra’s algorithm (infinite distances except the starting cell), iteratively updating route costs via priority queue simulation, and reconstructing the shortest route through parent-node backtracking.
4.2.3 AoT with Example (Algo-Shot) Prompt
@@
Algo-Direct: It integrates the AoT-Dijkstra prompt with the same example used in FewShot-Base - providing two cases in the grid, with each case including only the start point, end point, obstacles, and the optimal path computed using the Dijkstra algorithm.
Algo-Reasoning: Different from Algo-Direct, Algo-Reasoning combines AoT-Dijkstra with example that provides a detailed demonstration of the path planning reasoning steps from the starting point to the destination. To align with the reasoning process in this example, we configured the internal steps of AoT-Dijkstra in greater detail, clearly specifying how various situations encountered during path planning are handled.
@@ 5 Experimental Analysis
RQ1: How does the Algorithms of Thought (AoT) prompt perform compared to the Independent Route Planning prompt?
In Figure 3, we find that AoT prompting methods significantly outperform the Vanilla prompt across all metrics except AoT-DFS, especially in more complex environments. Moreover, the integration of LLMs with optimal pathfinding algorithms (e.g., Dijkstra or A*) achieves better planning performance compared to their integration with uninformed search methods like DFS. These methods show clear advantages in terms of FR and OR, indicating stronger path planning capabilities. However, a notable trade-off exists for AoT-DFS compared to Vanilla: despite the improvements in FR and OR, GM and MSE can sometimes be higher.
Compared to the more advanced independent prompting strategy CoT, AoT achieves comparable or even superior performance, particularly on Qwen2.5-7B, as illustrated in Figure 13. Figure 5 further highlights that AoT surpasses CoT in terms of geometric mean (GM), reflecting higher overall route quality. However, CoT exhibits more consistent performance across all models, with fewer failure cases as shown in Figure 2. In contrast, within the AoT framework, only AoT-Dijkstra consistently maintains robust results. These findings suggest that CoT’s step-by-step reasoning offers greater reliability across model sizes, while AoT’s effectiveness depends more on the model’s ability to accurately internalize and apply classical algorithmic logic.
FewShot-Base, despite its simplicity compared to advanced prompting strategies such as CoT and AoT, achieves performance comparable to AoT-Dijkstra by relying on minimal task demonstrations involving only obstacles, start/end points, and optimal paths. This highlights the efficacy of lightweight imitation-based learning and illustrates the strong transfer and generalization capabilities of LLMs in path planning tasks.
Overall, CoT demonstrates the lowest total error count, while AoT prompts shows comparable performance overall. However, AoT exhibits significant start/end mismatch issues in smaller models (e.g., Qwen2.5-7B).
@@
Our experiments comparing AoT-Dijkstra with different shot strategies (Figure 6) reveal that the Algo-Direct prompt generally enhances AoT-Dijkstra performance in most metrics, while the Algo-Reasoning setting improves its performance specifically in GM and MSE at the cost of reduced FR and OR. However, due to the limited capabilities of Qwen2.5-7B, Algo-Reasoning could not produce valid outputs on this smaller model. When using the same example, Algo-Direct prompting outperforms FewShot-Base by using the AoT-Dijkstra prompt while preserving example demonstration patterns.
@@
Figure 2 demonstrates that within the Qwen2.5 model series, scaling from 7B to 32B leads to substantial improvements in FR and RT. The 72B model further enhances FR while significantly boosting OR and reducing GM and MSE. This suggests that the FR gains observed in Qwen2.5-32B come at the cost of higher GM and MSE. Notably, Qwen2.5-72B achieves performance comparable to models such as GPT-4 Turbo(1 trillion) and DeepSeek-V3(671B), suggesting that simply increasing model size in GridRoute tasks eventually encounters a scaling law bottleneck. Additionally, as model size increases, errors such as illegal moves and start-end mismatches become significantly less frequent and are increasingly concentrated in path-through-obstacle cases. In most prompting strategies, inference speed tends to improve with larger model sizes, especially under AoT-Dijkstra and vanilla prompts. However, AoT-DFS shows less consistent improvement and even experiences relatively increased latency from 7B to 32B, indicating that response time is not strictly correlated with model size.
@@
As shown in Figure 7, the scaling of map size is a critical factor affecting the path planning performance across all evaluation metrics. As the map size increases, both the FR and OR exhibit a systematic decline, while the RT correspondingly increases. This scaling effect holds consistently regardless of model size, although the influence of model size diminishes as the map becomes larger—for instance, on a 30×30 grid, the performance of Qwen2.5-32B approaches that of Qwen2.5-72B. Notably, the nature of errors evolves with increasing map complexity, transitioning from primarily " Path though obstacles" to more intricate types of errors. While model size and architectural differences do influence performance, their impact is secondary to the fundamental constraints imposed by map dimensionality. In particular, the CoT prompting strategy exhibits more generalized adaptability across varying map sizes compared to the AoT approach. This suggests that with appropriate prompting strategies, LLMs can achieve path planning performance comparable to or even competitive with AoT methods.
@@ 6 Conclusion
We introduced GridRoute, a benchmark for systematically evaluating LLM-based route planning with guidance from classical algorithms. Our proposed AoT prompting framework enables LLMs to incorporate structured reasoning inspired by A*, Dijkstra, and DFS, significantly improving performance compared with Vallina, especially in complex environments. AoT performs comparably to or better than CoT, while example-based AoT variants further enhance generalization. Although CoT offers more stable performance across model sizes, our findings highlight the complementary strengths of symbolic algorithms and LLMs, suggesting promising directions for hybrid neuro-symbolic planning systems.
@@ A.1 Independent Route Planning Prompts
Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: -Right: -Left: -Down: -Up: 4. Build the path: - Start at ({start_x}, {start_y}). - At each step, evaluate valid moves and select the one that minimizes the distance to ({end_x}, {end_y}). 5. If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Example 1: Start: (3, 7), End: (4, 3), Obstacles: [((1, 2), (3, 4)), ((2, 5), (4, 6))] Path: [(3, 7), (4, 7), (5, 7), (5, 6), (5, 5), (5, 4), (5, 3), (4, 3)] Example 2: Start: (2, 4), End: (7, 5), Obstacles: [((3, 4), (5, 6))] Path: [(2, 4), (2, 3), (3, 3), (4, 3), (5, 3), (6, 3), (7, 3), (7, 4), (7, 5)] Return only the path as output, without additional explanations.
A.2 Algorithm of Thought(AoT) Prompts
@@
Plan a path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use DFS algorithm to explore possible paths. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves (same priority order): - Right: (x, y+1) - Left: (x, y-1) - Down: (x+1, y) - Up: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Depth-First Search (DFS) algorithm: - Initialize a stack with the start cell and its path [(start_x, start_y)] - Create a visited set to track explored cells - While stack is not empty: - Pop the last cell and its path from the stack - If current cell is the endpoint, return the complete path - If cell not in visited: - Mark cell as visited - Push all valid neighboring cells to the stack (FILO order): * For each neighbor in move order (Right→Left→Down→Up): - If neighbor is within grid bounds - If neighbor is not an obstacle - If neighbor not in visited - Push (neighbor_coords, path + [neighbor_coords]) to stack 5. Termination conditions: - If stack becomes empty before reaching endpoint, return empty path []. - First valid path found using DFS should be returned. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
@@
Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use A* algorithm to calculate the shortest path. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Right: (x, y+1) - Left: (x, y-1) - Down: (x+1, y) - Up: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use A* algorithm: - Initialize g_score (cost from start) for all cells to infinity, except the start cell which is 0. - Initialize f_score (g_score + heuristic) for all cells to infinity, except the start cell which is heuristic(start, end). - Use a priority queue to repeatedly select the grid cell with the lowest f_score. - For each selected cell, evaluate its valid neighbors: - Calculate the tentative g_score for the neighbor through the current cell. - If the tentative g_score is lower than the neighbor’s current g_score, update the neighbor’s parent, g_score, and f_score (g_score + heuristic(neighbor, end)). - Add the neighbor to the priority queue if it’s not already there or if this path is better. - Use the Manhattan distance as the heuristic function: h((x1, y1), (x2, y2)) = abs(x1 - x2) + abs(y1 - y2). - Continue until the endpoint (end_x, end_y) is reached or the priority queue is empty. 5. Construct the path: - Start at (start_x, start_y). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
@@
"Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use Dijkstra’s algorithm to calculate the shortest path. Follow these steps:" 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Right: (x, y+1) - Left: (x, y-1) - Down: (x+1, y) - Up: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Dijkstra’s algorithm: - Assign an initial distance of infinity to all grid cells except the starting point, which should have a distance of 0. - Use a priority queue to repeatedly select the grid cell with the smallest tentative distance. - For each selected cell, evaluate its valid neighbors and update their tentative distances if a shorter path is found. - Continue until the endpoint ({end_x}, {end_y}) is reached or all possible paths have been explored. 5. Construct the path:" - Start at ({start_x}, {start_y}). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
A.3 AoT with Example (Algo-Shot) Prompt

### v2

@@ Abstract
Recent advancements in Large Language Models (LLMs) have demonstrated their potential in planning and reasoning tasks, offering a flexible alternative to classical pathfinding algorithms. However, most existing studies focus on LLMs’ independent reasoning capabilities and overlook the potential synergy between LLMs and traditional algorithms. To fill this gap, we propose a comprehensive evaluation benchmark GridRoute to assess how LLMs can take advantage of traditional algorithms. We also propose a novel hybrid prompting technique called Algorithm of Planning (AoP), which introduces traditional algorithms’ guidance into prompting. Our benchmark evaluates six LLMs ranging from 7B to 72B parameters across various map sizes, assessing their performance in correctness, optimality, and efficiency in grid environments with varying sizes. Our results show that AoP significantly boosts performance across all model sizes, particularly in larger or more complex environments, suggesting a promising approach to addressing path planning challenges.111Our code is open-sourced at https://github.com/LinChance/GridRoute.
@@
Kechen Li1, Yaotian Tao2, Ximing Wen3, Quanwei Sun2, Zifei Gong4, Hongli Zhang 2, Chang Xu2, Xizhe Zhang2, Tianbo Ji1,††thanks: Corresponding author 1School of Transportation and Civil Engineering, Nantong University, China 2HongKong Linchance Intelligent Technology Co., Ltd. (LinChance), China 3College of Computing and Informatics, Drexel University, USA 4University of New South Wales, Australia kechen.li@stmail.ntu.edu.cn, {taoyaotian,quanweisun,HongliZhang,xuchang,zhangxizhe}@linchance.com xw384@drexel.edu, z3224950@zmail.unsw.edu.au, jitianbo@ntu.edu.cn
@@ 1 Introduction
To bridge this gap, we propose a comprehensive benchmark GridRoute to systematically evaluate the role of traditional algorithmic guidance in enhancing the pathfinding capabilities of LLMs. Our benchmark is designed around three core prompting paradigms: (1) independent prompting, (2) Algorithm of Planning(AoP) prompt, where traditional algorithms such as DFS, A*, and Dijkstra provide reasoning guidance; and (3) hybrid prompting, which leverages algorithm-generated trajectories or decisions within both direct-answer and reasoning-based prompts. This design facilitates a comprehensive comparison of traditional algorithmic guidance versus independent prompting approaches, and further allows for in-depth analysis of variation within algorithm-guided prompts.
@@
Through this benchmark, we aim to answer several key questions: To what extent can LLMs benefit from classical search heuristics? How do the different shot strategies affect the traditional algorithm’s guidance for LLM? Do algorithm-guided prompt performance remain indispensable regardless of model scale and map size? Our findings suggest that algorithm-guided prompts achieve better results than vanilla prompts and comparable performance compared with the CoT prompt. While larger models tend to perform better under independent prompting, algorithm-guided strategies significantly boost performance across all model sizes, particularly in larger or more complex environments. Moreover, integrating few-shot learning could further improve the performance of algorithm-guided prompt. These results highlight the complementary strengths of LLMs and classical algorithms and point toward promising directions for hybrid neuro-symbolic planning systems.
@@
We introduce AoP prompting, a novel framework that guides LLMs using classical pathfinding algorithms (e.g., A*, Dijkstra, DFS) by embedding their reasoning principles directly into the prompt.
We perform large-scale experiments across multiple LLMs and environment sizes, showing that AoP consistently improves planning performance, especially in complex settings, and offers complementary benefits to traditional CoT prompting.
@@ 4.2.1 Independent Route Planning Prompt
4.2.2 Algorithms of Planning (AoP) Prompt
To evaluate the capability of LLMs in applying classical algorithms, we designed the Algorithms of Planning (AoP) prompting framework. This framework guides LLMs to mimic classical path planning algorithms by prompting them to apply algorithmic reasoning, analyze the task environment, constrain movement directions, explain algorithmic principles, and adhere to a clear output format.
AoP-DFS: This prompt is designed to guide the model in following the Depth-First Search (DFS).Instead of outlining the algorithm formally, the prompt incorporates its core mechanisms—ordered expansion, backtracking, and visit marking—into the structure of the instruction itself. By structuring the order of exploration and discouraging revisits,it leads the model to follow a behavior pattern similar to DFS.
AoP-A*: This prompt adapts the logic of the A* algorithm to guide the model in path planning. It introduces key elements of A*, such as cost estimation, heuristic evaluation, and goal-directed prioritization. The prompt defines valid moves, explains how to measure distance using the Manhattan heuristic, and describes how to compare path options. The model is prompted to evaluate each possible move based on its distance to the goal and the cost incurred so far, simulating how A* selects efficient paths.
AoP-Dijkstra: It provides a structured framework by embedding Dijkstra’s core logic – greedy distance minimization, neighbor evaluation, and path optimality. In detail, it enforces a multi-step workflow: initializing Dijkstra’s algorithm (infinite distances except the starting cell), iteratively updating route costs via priority queue simulation, and reconstructing the shortest route through parent-node backtracking.
4.2.3 AoP with Example (Algo-Shot) Prompt
@@
Algo-Direct: It integrates the AoP-Dijkstra prompt with the same example used in FewShot-Base - providing two cases in the grid, with each case including only the start point, end point, obstacles, and the optimal path computed using the Dijkstra algorithm.
Algo-Reasoning: Different from Algo-Direct, Algo-Reasoning combines AoP-Dijkstra with example that provides a detailed demonstration of the path planning reasoning steps from the starting point to the destination. To align with the reasoning process in this example, we configured the internal steps of AoP-Dijkstra in greater detail, clearly specifying how various situations encountered during path planning are handled.
@@ 5 Experimental Analysis
RQ1: How does the Algorithms of Planning (AoP) prompt perform compared to the Independent Route Planning prompt?
In Figure 3, we find that AoP prompting methods significantly outperform the Vanilla prompt across all metrics except AoP-DFS, especially in more complex environments. Moreover, the integration of LLMs with optimal pathfinding algorithms (e.g., Dijkstra or A*) achieves better planning performance compared to their integration with uninformed search methods like DFS. These methods show clear advantages in terms of FR and OR, indicating stronger path planning capabilities. However, a notable trade-off exists for AoP-DFS compared to Vanilla: despite the improvements in FR and OR, GM and MSE can sometimes be higher.
Compared to the more advanced independent prompting strategy CoT, AoP achieves comparable or even superior performance, particularly on Qwen2.5-7B, as illustrated in Figure 12.
However, CoT exhibits more consistent performance across all models, with fewer failure cases as shown in Figure 2. In contrast, within the AoP framework, only AoP-Dijkstra consistently maintains robust results. These findings suggest that CoT’s step-by-step reasoning offers greater reliability across model sizes, while AoP’s effectiveness depends more on the model’s ability to accurately internalize and apply classical algorithmic logic.
FewShot-Base, despite its simplicity compared to advanced prompting strategies such as CoT and AoP, achieves performance comparable to AoP-Dijkstra by relying on minimal task demonstrations involving only obstacles, start/end points, and optimal paths. This highlights the efficacy of lightweight imitation-based learning and illustrates the strong transfer and generalization capabilities of LLMs in path planning tasks.
Overall, CoT demonstrates the lowest total error count, while AoP prompts shows comparable performance overall. However, AoP exhibits significant start/end mismatch issues in smaller models (e.g., Qwen2.5-7B).
@@
Our experiments comparing AoP-Dijkstra with different shot strategies (Figure 5) reveal that the Algo-Direct prompt generally enhances AoP-Dijkstra performance in most metrics, while the Algo-Reasoning setting improves its performance specifically in GM and MSE at the cost of reduced FR and OR. However, due to the limited capabilities of Qwen2.5-7B, Algo-Reasoning could not produce valid outputs on this smaller model. When using the same example, Algo-Direct prompting outperforms FewShot-Base by using the AoP-Dijkstra prompt while preserving example demonstration patterns.
@@
Figure 2 demonstrates that within the Qwen2.5 model series, scaling from 7B to 32B leads to substantial improvements in FR and RT. The 72B model further enhances FR while significantly boosting OR and reducing GM and MSE. This suggests that the FR gains observed in Qwen2.5-32B come at the cost of higher GM and MSE. Notably, Qwen2.5-72B achieves performance comparable to models such as GPT-4 Turbo(1 trillion) and DeepSeek-V3(671B), suggesting that simply increasing model size in GridRoute tasks eventually encounters a scaling law bottleneck. Additionally, as model size increases, errors such as illegal moves and start-end mismatches become significantly less frequent and are increasingly concentrated in path-through-obstacle cases. In most prompting strategies, inference speed tends to improve with larger model sizes, especially under AoP-Dijkstra and vanilla prompts. However, AoP-DFS shows less consistent improvement and even experiences relatively increased latency from 7B to 32B, indicating that response time is not strictly correlated with model size.
@@
As shown in Figure 6, the scaling of map size is a critical factor affecting the path planning performance across all evaluation metrics. As the map size increases, both the FR and OR exhibit a systematic decline, while the RT correspondingly increases. This scaling effect holds consistently regardless of model size, although the influence of model size diminishes as the map becomes larger—for instance, on a 30×30 grid, the performance of Qwen2.5-32B approaches that of Qwen2.5-72B. Notably, the nature of errors evolves with increasing map complexity, transitioning from primarily " Path though obstacles" to more intricate types of errors. While model size and architectural differences do influence performance, their impact is secondary to the fundamental constraints imposed by map dimensionality. In particular, the CoT prompting strategy exhibits more generalized adaptability across varying map sizes compared to the AoP approach. This suggests that with appropriate prompting strategies, LLMs can achieve path planning performance comparable to or even competitive with AoP methods.
@@ 6 Conclusion
We introduced GridRoute, a benchmark for systematically evaluating LLM-based route planning with guidance from classical algorithms. Our proposed AoP prompting framework enables LLMs to incorporate structured reasoning inspired by A*, Dijkstra, and DFS, significantly improving performance compared with Vallina, especially in complex environments. AoP performs comparably to or better than CoT, while example-based AoP variants further enhance generalization. Although CoT offers more stable performance across model sizes, our findings highlight the complementary strengths of symbolic algorithms and LLMs, suggesting promising directions for hybrid neuro-symbolic planning systems.
@@ A.1 Independent Route Planning Prompts
Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: -Up: -Down: -Right: -Left: 4. Build the path: - Start at ({start_x}, {start_y}). - At each step, evaluate valid moves and select the one that minimizes the distance to ({end_x}, {end_y}). 5. If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Example 1: Start: (3, 7), End: (4, 3), Obstacles: [((1, 2), (3, 4)), ((2, 5), (4, 6))] Path: [(3, 7), (4, 7), (5, 7), (5, 6), (5, 5), (5, 4), (5, 3), (4, 3)] Example 2: Start: (2, 4), End: (7, 5), Obstacles: [((3, 4), (5, 6))] Path: [(2, 4), (2, 3), (3, 3), (4, 3), (5, 3), (6, 3), (7, 3), (7, 4), (7, 5)] Return only the path as output, without additional explanations.
A.2 Algorithm of Planning(AoP) Prompts
@@
Plan a path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use DFS algorithm to explore possible paths. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves (same priority order): - Up: (x, y+1) - Down: (x, y-1) - Right: (x+1, y) - Left: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Depth-First Search (DFS) algorithm: - Initialize a stack with the start cell and its path [(start_x, start_y)] - Create a visited set to track explored cells - While stack is not empty: - Pop the last cell and its path from the stack - If current cell is the endpoint, return the complete path - If cell not in visited: - Mark cell as visited - Push all valid neighboring cells to the stack (FILO order): * For each neighbor in move order (Right→Left→Down→Up): - If neighbor is within grid bounds - If neighbor is not an obstacle - If neighbor not in visited - Push (neighbor_coords, path + [neighbor_coords]) to stack 5. Termination conditions: - If stack becomes empty before reaching endpoint, return empty path []. - First valid path found using DFS should be returned. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
@@
Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use A* algorithm to calculate the shortest path. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Up: (x, y+1) - Down: (x, y-1) - Right: (x+1, y) - Left: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use A* algorithm: - Initialize g_score (cost from start) for all cells to infinity, except the start cell which is 0. - Initialize f_score (g_score + heuristic) for all cells to infinity, except the start cell which is heuristic(start, end). - Use a priority queue to repeatedly select the grid cell with the lowest f_score. - For each selected cell, evaluate its valid neighbors: - Calculate the tentative g_score for the neighbor through the current cell. - If the tentative g_score is lower than the neighbor’s current g_score, update the neighbor’s parent, g_score, and f_score (g_score + heuristic(neighbor, end)). - Add the neighbor to the priority queue if it’s not already there or if this path is better. - Use the Manhattan distance as the heuristic function: h((x1, y1), (x2, y2)) = abs(x1 - x2) + abs(y1 - y2). - Continue until the endpoint (end_x, end_y) is reached or the priority queue is empty. 5. Construct the path: - Start at (start_x, start_y). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
@@
Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use Dijkstra’s algorithm to calculate the shortest path. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Up: (x, y+1) - Down: (x, y-1) - Right: (x+1, y) - Left: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Dijkstra’s algorithm: - Assign an initial distance of infinity to all grid cells except the starting point, which should have a distance of 0. - Use a priority queue to repeatedly select the grid cell with the smallest tentative distance. - For each selected cell, evaluate its valid neighbors and update their tentative distances if a shorter path is found. - Continue until the endpoint ({end_x}, {end_y}) is reached or all possible paths have been explored. 5. Construct the path:" - Start at ({start_x}, {start_y}). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
A.3 AoP with Example (Algo-Shot) Prompt